OpenStudy (wasiqss):

find length of the curve {(x/2)-(x^2/4)+(ln(1-x)/2} between x=0 and x= .5

myininaya (myininaya):

\[\int\limits_{0}^{\frac{1}{2}}\sqrt{1+(y')^2} dx\]

myininaya (myininaya):

so have you found y' yet?

myininaya (myininaya):

\[y'=\frac{1}{2}-\frac{1}{2}x+\frac{1}{2} \cdot \frac{-1}{1-x}\] you don't have enough closed parenthesis so i don't know if i read that right

OpenStudy (wasiqss):

i ll draw it

OpenStudy (wasiqss):